OpenStudy (anonymous):

Solve 4x2 = 12x + 40.

OpenStudy (anonymous):

4x^2 - 12x + 40 = 0 4(x^2 - 3x + 10) = 0 x^2 - 3x + 10 = 0

OpenStudy (anonymous):

move all the terms to the left sides: 4x^2-12x-40=0 Now notice we can factor out a 4, resulting in: 4(x^2-3x-10) Now you have to find factor just the part inside the parenthese, to do that find two numbers that multiply to be -10 and add to be -3

OpenStudy (anonymous):

4x^2-12x -40 =0 x=-2 x=5